OpenStudy (anonymous):

When the integral g(x) from -2 to 5 is equal to 5. What is the integral from -2 to 5 of (g(x)+2)dx=? I thought it would be 7, but the teacher as an answer of 19. Please explain thought process.

hartnn (hartnn):

the integral, gets distributed to both terms. \[\int g(x)dx+\int 2 dx\] so, you need to solve 2nd integral separately, [1st integral=5] can you ?

OpenStudy (anonymous):

i just did the same, but i didn't get 19...

OpenStudy (anonymous):

@hartnn i did that but got 11. unless the -2 becomes a positive 2 somehow, i don't see how you get 19.

hartnn (hartnn):

5+2(5-(-2)) = +19

OpenStudy (phi):

what is \[2\int\limits_{-2}^{5}dx\]

OpenStudy (anonymous):

YUP! come on and say int dude: "F that teacher!"

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

6

OpenStudy (anonymous):

The teacher is right though?

OpenStudy (whpalmer4):

If you think of the integral definition, it's just summing up the little rectangles under the curve. In the attached figure, the light blue area is g(x) integrated from x = -2 to x = 5, and has area = l*w = 7*(5/7) = 5. When we switch to integrating g(x) + 2 over that same portion of the x axis, we add in the green region, and now our area is increased by the area l*w = 2*7 = 14, giving a total of 5+14 = 19.

OpenStudy (whpalmer4):

Sorry, bad label in my graph, g(x) = 5/7, not 5x/7!

OpenStudy (zehanz):

@IStutts : although teachers sometimes make mistakes (I know, because I am a teacher), in this case the teacher's right, just read @whpalmer4's answer...

OpenStudy (anonymous):

hartnn's description makes the most sense to me mathmatically. Thanks for the help!! :)

hartnn (hartnn):

welcome ^_^